OpenStudy (anonymous):

Describe how to check the product of two rational expressions. I have the answer, but how do I check if I'm right?

OpenStudy (lopus):

\[\frac{ 2 }{ 3 }*\frac{ 4 }{ 2 } =\frac{ 8 }{ 6 }\]

OpenStudy (lopus):

\[\frac{ a }{ b }*\frac{ b }{ c }=\frac{ ab}{ bc }\]

OpenStudy (anonymous):

Thank you :)